A pulsed laser has an average output power of 12W. Each pulse consists of light at
wavelength 400nm and has 16ms duration. Compute the number of photons that are emitted
in a single pulse?
1
Expert's answer
2020-06-09T13:19:39-0400

Electrical energy of the laser:


"E=Pt."

Energy o the emitted photons:


"E=n\frac{hc}{\lambda}."

Hence:


"Pt=n\frac{hc}{\lambda},\\\space\\
n=\frac{Pt\lambda}{hc}=3.866\cdot10^{17}."